I honestly do not know where to begin on this one. I looked in the textbook for an example and could not find one for this setup.
Word problem in a form of a Linear equation.
PROBLEM: For a summer job, Dwayne decides to sell magazine subscriptions. He will be paid $18 a day plus $1 for each subscription he sells. If he works for 25 days and makes $574, how many subscriptions did he sell?

Answer by richwmiller(17219)


18d+1s=p
18*25+1s=574
9*50+s=574
450+s=574
s=124
He sold 124 subscriptions in 25 days